What volume of O2 would be required to burn 224 mL of H2 at 273K and 1.013 bar?

2H2 + O2 → 2H2O


1
Expert's answer
2021-05-17T04:20:11-0400

273 K and 1.013 bar are STP conditions, so 1 mol of gas occupies volume of 22.4 L

n(H2) "= \frac{0.224}{22.4} = 0.01 \;mol"

According to the reaction:

n(O2) "= \frac{1}{2}n(H_2) = \frac{1}{2} \times 0.01 = 0.005 \;mol"

V(O2) "= 0.005 \times 22.4 = 0.112 \;L = 112 \;mL"

Answer: 112 mL